if a particle of mass m in a horizontal circle of radius r with a centripetal force(-k/r^2) the total energy is (1)-k/2r (2)-k/r (3)-2k/r (4)-4k/r

Grade:11

1 Answers

sanatan sharma
11 years ago

centripetal force (F) = - k/r2    

F = - dU/dr      ( U = potential energy )

kdr/r2 = dU

integratin    we get    U = -k/r

  F = mv2/r = k/r2     ( -ve sign neglected bcoz it just indicates the direction )

  mv2 = k/r

kinetic energy = 0.5mv2 = k/2r 

total energy = KE + PE = -k/2r

 

PLEASE HIT LIKE IF YOU LIKE THE SOLN